A committee of four is chosen at random from a group of 8 women and 5 men. Find the probability that the committee contains at least one man.
Answer by jim_thompson5910(35256)

There are 8 C 4 = 70 ways to choose a group with all women.

Note: I'm using the formula shown here

There are 13 C 4 = 715 ways to choose a group of men and women.

So there are 715 - 70 = 645 ways to choose a group with at least one man.

This means the probability is 645/715 = 129/143 = 0.9020979 (so roughly a 90.20979% chance)